OpenStudy (anonymous):

The sum of two numbers is 170 and their difference is 26. Find the two numbers. Show how to solve this using any method.

OpenStudy (anonymous):

Let the two nos be x and y x+y= 170 x-y=26 cross out the y 2x=196 x=98 y= 170-98 y= 72

OpenStudy (anonymous):

The sides of a triangle have measures of x, x, and y, and the perimeter of the triangle is 46. The measure y is 4 centimeters longer than the measure x. Write the solution for this as an ordered pair.